# Malle’s conjecture and Brauer groups of stacks

### v2

@@ 9.1. Fair heights
Assume that $H$ is fair. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
We also give a version of the conjecture for field extensions. In the statement we consider field extensions up to isomorphism, denote by $\widetilde{K}$ the Galois closure of $K$, and consider $\operatorname{Gal}(\widetilde{K}/k)$ as a permutation group. This is the original setting for Malle’s conjecture, and we give a precise prediction for the asymptotic providing the discriminant is fair.
@@
Let $G\subseteq S_{n}$ be a transitive subgroup with normaliser $N$ and centraliser $C$. Assume that the discriminant is a fair height function. Then
@@
9.2. Unfair heights
Assume now that $H$ is not necessarily fair. Here one reduces to the fair case as follows. Let $M:=\langle\mathcal{M}(L)\rangle\subseteq G$ be the subgroup scheme generated by the minimal weight conjugacy classes with respect to $H$. We consider the associated (adjoint) Iitaka fibration $BG\to B(G/M)$ from §3.5. The counting problem is then given as a sum of the counting problems of the fibres of the Iitaka fibration, with the restriction of $H$ to each fibre now being fair by Lemma 3.31. So the setting from §9.1 applies. We thus apply Conjecture 9.1 to each fibre, and keep track of groupoid cardinalities using Lemmas 2.9 and 2.14(1), with the expectation that the sum over all the fibres converges. This leads to the following.
@@ 9.2. Unfair heights
Here $M_{\psi}$ denotes the inner twist of $M$ by a lift of $\psi$ along $BG(k)\to B(G/M)(k)$, as defined in Definition 2.12; this is a minor abuse of notation which naturally requires that $\psi$ is in the image of $BG(k)$, and $BM_{\psi}$ is independent of the choice of lift. The stack $BM_{\psi}$ is a fibre of the Iitaka fibration by Lemma 2.13. The constant $c(k,M_{\psi},H)$ is as in Conjecture 9.1; this is admissible as the pull-back of $H$ to $BM_{\psi}$ is fair by Lemma 3.31. If $G$ is constant, this fibre is exactly the collection of $G$-extensions of $k$ which realise a given $G/M$-extension.
@@
Warning: even if $G$ is constant, the group scheme $M_{\psi}$ need not be constant (this happens when counting $D_{4}$-quartics of bounded discriminant; see §10.2). So in Malle’s conjecture one needs to work with general height functions on finite étale group schemes, even if one is only interested in number fields of bounded discriminant. The appearance of finite étale group schemes in Malle’s conjecture appears to have been first noticed by Tükelli [87, Conj. 4.8] (see also [4, Lem. 1.3] and [33, §9.5]).
@@ 9.3. The total count
We explain this process in detail when $G$ is a finite group and $\Omega$ is the thin set from Theorem 3.40(2). We represent elements of $BG(k)$ by homorphisms $\Gamma_{k}\to G$ of bounded height $H$. Writing the thin set from Theorem 3.40(2) as a union of images of $BT\to BG$ corresponds to stratifying all homorphisms $\Gamma_{k}\to G$ according to whether they lift suitable cyclotomic fields; this latter perspective is also considered in a recent paper of Wang [89]. The leading constant obtained will be quite complicated, as it can be a mix of both fair and unfair cases, and finite étale group schemes (even though $G$ itself is constant). This leads to
@@
Klüners [50] was the first to observe that fields which contain a given cyclotomic field can be accumulating in Malle’s conjecture. Türkelli [87, Conj 4.8] proposed a possible fix to Malle’s conjecture by taking into account these cyclotomic fields. A recent paper of Wang [89, Conj. 7] gives a counter-example to Türkelli’s original conjecture and suggests a modification for finding the power of $\log B$, with a key observation that one discounts those subfields of $k(\mu_{\exp(G)})$ for which no lift exists. A priori it could be possible that one should remove further thin sets from each $BN_{\varphi_{\psi}}$; but we have shown in Lemma 9.7 that this is unnecessary. Our work builds on Wang’s paper by also putting forward a conjecture for the leading constant.
@@ 9.4. Equidistribution
In the Manin’s conjecture literature the property is called equidistribution, since it stipulates weak convergence of a sequence of measures, and was first considered by Peyre [71, §3] (see [73, 3.21] for this perspective, as well as the Portmanteau theorem [59, Thm. 13.16] for various equivalent formulations). With the relevant Tamagawa measure now in hand for $BG$, we are able to state the following equidistribution conjecture. Recall that a continuity set is set whose boundary has measure 00.
@@
Assume that $H$ is fair. Let $W\subseteq\prod_{v}BG[k_{v}]$ be a continuity set. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
Peyre shows in [71, Prop. 3.3] in the setting of Manin’s conjecture that the equidistribution property is equivalent to proving his asymptotic with respect to arbitrary choices of adelic metric on the line bundle (see also [26, Prop. 2.10] for a general topological statement). We have an analogous property in our case.
@@
The equidistribution conjecture adds further evidence to the need to remove a thin set, since in general equidistribution can only hold after removal of a thin set (this has been shown for example in the case of cubic extensions of bounded radical discriminant [82, Thm. 6]). But more than that, once equidistribution has been obtained, the following result shows that removing an additional thin set does not change the asymptotic formula. This means that actually there is no danger in removing a larger thin set than required in Conjecture 9.1. This property for Manin’s conjecture was first verified in [22, Thm. 1.2].
@@ 9.5. Strong equidistribution
Assume that $H$ is fair. Let $W\subseteq BG[\mathbf{A}_{k}]_{\mathcal{M}(L)}$ be a continuity set. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
Peyre [71, §3] only considered equidistribution for anticanonical heights. We could not find any results in the Manin’s conjecture literature regarding equidistribution for non-anticanonical heights. However it seems very reasonable to expect that the equidistribution property should hold for height functions associated to big adjoint rigid line bundles, providing equidistribution takes place in the adelic points of the open subset given by removing the adjoint divisor. Conjecture 9.15 is exactly modelled on this situation (cf. Remark 3.17).
9.6. Equidistribution for unfair heights
If $H$ is not fair then we follow the procedure from §9.2. Namely we pass to the Iitaka fibration, and then Conjecture 9.10 predicts that the rational points in each fibre are equidistributed with respect to the induced Tamagawa measure on the fibre.
This explains the phenomenon, first observed by Wood [91, Prop. 1.4] when counting abelian extensions of bounded discriminant, that the equidistribution property need not hold when counting all extensions. The fix is very simple: to get equidistribution one should only count extensions which lie in a given fibre of the Iitaka fibration.
@@ 9.6. Equidistribution for unfair heights
Darda and Yasuda define in [32, §3.5] a Radon measure in the case of abelian $G$ via an abstract process. It is not clear how this relates to our Tamagawa measure. The fact that they obtain equidistribution in [32, Thm. 3.5.8] for unfair heights suggests it is of a very different nature.
@@ 9.7. Products
We next consider compatibility of our conjectures with respect to products. We focus on the fair case as it is more fundamental. Let $G_{1}$ and $G_{2}$ be finite étale group schemes over $k$ with fair heights $H_{1}$ and $H_{2}$. We define a height $H:=H_{1}\boxtimes H_{2}$ on $G:=G_{1}\times G_{2}$ by taking the product of height functions. To ensure that a product of fair heights is fair, we assume that the Fujita invariants are equal.
@@ 10.1. $S_{n}$-extensions
The orbifold line bundle $\Delta$ is fair on $BS_{n}$.
@@
Therefore in this case the adjoint Iitaka fibration is trivial and the discriminant is a fair height function, thus we are in the situation of Conjecture 9.1.
@@ 10.2.2. Discriminant
The subgroup generated by the minimal index elements is $M=\{(),(1,3),(2,4),(1,3)(2,4)\}$. The adjoint Iitaka fibration of the orbifold line bundle $\Delta$ is given by $BD_{4}\to BC_{2}$. In particular $\Delta$ is not fair.
@@
The discriminant is not a fair height function in this case. The adjoint Iitaka fibration is by definition the map which sends a $D_{4}$-extension $L/k$ to its quadratic subfield $L^{(1,3),(2,4)}$. By the Galois correspondence this is the unique quadratic subfield of the corresponding $D_{4}$-quartic $L^{(1,3)}$.
@@ 10.2.3. An Artin conductor
We have $a(C)=1,b(C)=2$. The orbifold line bundle $C$ is fair.
@@ 10.3. Klüners’s counter-example
The group generated by these two conjugacy classes is $C_{3}\times C_{3}$ so the discriminant is not fair and the Iitaka fibration is $BG\to BC_{2}$.
@@ 10.4. A new dihedral counter-example
In the case $n=3$ we obtain $S_{3}$, and one is counting non-cyclic cubic fields of bounded radical discriminant. The collection of points in the image of $BG_{3}$ is weakly accumulating: it needs to be removed before equidistribution is obtained. These correspond to exactly the pure cubic extensions; cf. Example 3.37. In fact this has been proven by Shankar and Thorne [82, Thm. 4].
@@ 10.5. $A_{4}$-quartics of bounded discriminant
$2\#\left\{[K:\mathbb{Q}]=4:|\Delta_{K}|\leq B,\operatorname{Gal}(\widetilde{K}/\mathbb{Q})\cong A_{4}\right\}\sim c(\mathbb{Q},A_{4},D)B^{1/2}\log B,$
@@
$c(\mathbb{Q},A_{4},D)=\frac{35}{648}\prod_{p>3}\left(1-\frac{1}{p}\right)^{2}\left(1+\frac{2+\left(\frac{-3}{p}\right)}{p}\right).$
There is in fact already a conjecture regarding this problem in [29, §2.7], but the expression is more complicated than ours and involves a limit process and sorting $A_{4}$-quartics by their cubic resolvent. It is not immediate how to compare the two conjectures directly. Nonetheless, in [29, §2.7] the numerical value $0.074\ldots$ is given for the leading constant, and a simple computation shows that our constant satisfies $2c(\mathbb{Q},A_{4},D)=0.0729\ldots$ after considering the product over $0003<p<100,$ 000. (In private communication, Henri Cohen has confirmed that this minor discrepancy is not a problem and an artifact of using a least squares approximation in their calculation). The convergence is very slow, as one would expect, since it involves a conditionally convergent product; a faster converging product can be obtained by using the convergence factors from Theorem 8.17. We count isomorphism classes of fields, whereas [29] counts subfields of $\bar{\mathbb{Q}}$. To compare the two counts, we note that there are $4$ distinct embeddings of any $A_{4}$-quartic into $\bar{\mathbb{Q}}$, which matches with our factor $2$ in front and additional factor $1/2$ in the leading constant.
@@
with the latter two classes being swapped by $\operatorname{Gal}(\mathbb{Q}(\zeta_{3})/\mathbb{Q})$. It follows that $a(D)=1/2$ and $b(\mathbb{Q},D)=2$ (taking into account the Galois action). Since all conjugacy classes have minimal index, it follows that together they generate the group and hence the Iitaka fibration is trivial. Moreover the discriminant is the square of an anticanonical height; this may explain why this problem is so difficult, since the anticanonical height is generally the most difficult height in Manin’s conjecture.
We have $\widehat{A}_{4}=\mu_{3}$, so there are no non-trivial $1$-dimensional characters defined over $\mathbb{Q}$. The effective cone constant is thus $a(L)^{b(\mathbb{Q},L)-1}=1/2$. Since our height is a multiple of an anticanonical height, the relevant Brauer group is the unramified Brauer group.
@@
Finally Conjecture 9.3 states that we should remove $\widetilde{K}$ for which $\widetilde{K}\cap\mathbb{Q}(\mu_{12})\neq\mathbb{Q}$. However there are no such fields; indeed we have $\mathbb{Q}(\mu_{12})=\mathbb{Q}(i,\sqrt{3})$, but $\widetilde{K}$ cannot contain a quadratic subfield as $A_{4}$ has no subgroup of index $2$.
@@ 10.6. $A_{4}$-quartics with a transcendental Brauer group
We now consider a different height function which exhibits new behaviour. Namely, there exists a fair height function on $A_{4}$ for which the partially unramified transcendental Brauer group is non-trivial.
@@ 10.6.1. Numerical experiments
Let $\mathcal{C}$ consist of the minimal conjugacy classes $(1,2,3),(1,3,2)$. Note that we have $a(C)=1$ and $b(\mathbb{Q},C)=1$. The height is fair since $(1,2,3)$ and its conjugate $(1,2,4)$ generate $A_{4}$.

### v3

@@ 9.1. Balanced heights
Assume that $H$ is balanced. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
We also give a version of the conjecture for field extensions. In the statement we consider field extensions up to isomorphism, denote by $\widetilde{K}$ the Galois closure of $K$, and consider $\operatorname{Gal}(\widetilde{K}/k)$ as a permutation group. This is the original setting for Malle’s conjecture, and we give a precise prediction for the asymptotic providing the discriminant is balanced.
@@
Let $G\subseteq S_{n}$ be a transitive subgroup with normaliser $N$ and centraliser $C$. Assume that the discriminant is a balanced height function. Then
@@
9.2. Unbalanced heights
Assume now that $H$ is not necessarily balanced. Here one reduces to the balanced case as follows. Let $M:=\langle\mathcal{M}(L)\rangle\subseteq G$ be the subgroup scheme generated by the minimal weight conjugacy classes with respect to $H$. We consider the associated (adjoint) Iitaka fibration $BG\to B(G/M)$ from §3.5. The counting problem is then given as a sum of the counting problems of the fibres of the Iitaka fibration, with the restriction of $H$ to each fibre now being balanced by Lemma 3.31. So the setting from §9.1 applies. We thus apply Conjecture 9.1 to each fibre, and keep track of groupoid cardinalities using Lemmas 2.9 and 2.14(1), with the expectation that the sum over all the fibres converges. This leads to the following.
@@ 9.2. Unbalanced heights
Here $M_{\psi}$ denotes the inner twist of $M$ by a lift of $\psi$ along $BG(k)\to B(G/M)(k)$, as defined in Definition 2.12; this is a minor abuse of notation which naturally requires that $\psi$ is in the image of $BG(k)$, and $BM_{\psi}$ is independent of the choice of lift. The stack $BM_{\psi}$ is a fibre of the Iitaka fibration by Lemma 2.13. The constant $c(k,M_{\psi},H)$ is as in Conjecture 9.1; this is admissible as the pull-back of $H$ to $BM_{\psi}$ is balanced by Lemma 3.31. If $G$ is constant, this fibre is exactly the collection of $G$-extensions of $k$ which realise a given $G/M$-extension.
@@
Warning: even if $G$ is constant, the group scheme $M_{\psi}$ need not be constant (this happens when counting $D_{4}$-quartics of bounded discriminant; see §10.2). So in Malle’s conjecture one needs to work with general height functions on finite étale group schemes, even if one is only interested in number fields of bounded discriminant. The appearance of finite étale group schemes in Malle’s conjecture appears to have been first noticed by Tükelli [89, Conj. 4.8] (see also [4, Lem. 1.3] and [33, §9.5]).
@@ 9.3. The total count
We explain this process in detail when $G$ is a finite group and $\Omega$ is the thin set from Theorem 3.40(2). We represent elements of $BG(k)$ by homorphisms $\Gamma_{k}\to G$ of bounded height $H$. Writing the thin set from Theorem 3.40(2) as a union of images of $BT\to BG$ corresponds to stratifying all homorphisms $\Gamma_{k}\to G$ according to whether they lift suitable cyclotomic fields; this latter perspective is also considered in a recent paper of Wang [91]. The leading constant obtained will be quite complicated, as it can be a mix of both balanced and unbalanced cases, and finite étale group schemes (even though $G$ itself is constant). This leads to
@@
Klüners [50] was the first to observe that fields which contain a given cyclotomic field can be accumulating in Malle’s conjecture. Türkelli [89, Conj 4.8] proposed a possible fix to Malle’s conjecture by taking into account these cyclotomic fields. A recent paper of Wang [91, Conj. 7] gives a counter-example to Türkelli’s original conjecture and suggests a modification for finding the power of $\log B$, with a key observation that one discounts those subfields of $k(\mu_{\exp(G)})$ for which no lift exists. A priori it could be possible that one should remove further thin sets from each $BN_{\varphi_{\psi}}$; but we have shown in Lemma 9.7 that this is unnecessary. Our work builds on Wang’s paper by also putting forward a conjecture for the leading constant.
@@ 9.4. Equidistribution
In the Manin’s conjecture literature the property is called equidistribution, since it stipulates weak convergence of a sequence of measures, and was first considered by Peyre [73, §3] (see [75, 3.21] for this perspective, as well as the Portmanteau theorem [60, Thm. 13.16] for various equivalent formulations). With the relevant Tamagawa measure now in hand for $BG$, we are able to state the following equidistribution conjecture. Recall that a continuity set is set whose boundary has measure 00.
@@
Assume that $H$ is balanced. Let $W\subseteq\prod_{v}BG[k_{v}]$ be a continuity set. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
Peyre shows in [73, Prop. 3.3] in the setting of Manin’s conjecture that the equidistribution property is equivalent to proving his asymptotic with respect to arbitrary choices of adelic metric on the line bundle (see also [26, Prop. 2.10] for a general topological statement). We have an analogous property in our case.
@@
The equidistribution conjecture adds further evidence to the need to remove a thin set, since in general equidistribution can only hold after removal of a thin set (this has been shown for example in the case of cubic extensions of bounded radical discriminant [84, Thm. 6]). But more than that, once equidistribution has been obtained, the following result shows that removing an additional thin set does not change the asymptotic formula. This means that actually there is no danger in removing a larger thin set than required in Conjecture 9.1. This property for Manin’s conjecture was first verified in [22, Thm. 1.2].
@@ 9.5. Strong equidistribution
Assume that $H$ is balanced. Let $W\subseteq BG[\mathbf{A}_{k}]_{\mathcal{M}(L)}$ be a continuity set. Then there exists a thin subset $\Omega\subseteq BG[k]$ such that
@@
Peyre [73, §3] only considered equidistribution for anticanonical heights. We could not find any results in the Manin’s conjecture literature regarding equidistribution for non-anticanonical heights. However it seems very reasonable to expect that the equidistribution property should hold for height functions associated to big adjoint rigid line bundles, providing equidistribution takes place in the adelic points of the open subset given by removing the adjoint divisor. Conjecture 9.15 is exactly modelled on this situation (cf. Remark 3.17).
9.6. Equidistribution for unbalanced heights
If $H$ is not balanced then we follow the procedure from §9.2. Namely we pass to the Iitaka fibration, and then Conjecture 9.10 predicts that the rational points in each fibre are equidistributed with respect to the induced Tamagawa measure on the fibre.
This explains the phenomenon, first observed by Wood [93, Prop. 1.4] when counting abelian extensions of bounded discriminant, that the equidistribution property need not hold when counting all extensions. The fix is very simple: to get equidistribution one should only count extensions which lie in a given fibre of the Iitaka fibration.
@@ 9.6. Equidistribution for unbalanced heights
Darda and Yasuda define in [32, §3.5] a Radon measure in the case of abelian $G$ via an abstract process. It is not clear how this relates to our Tamagawa measure. The fact that they obtain equidistribution in [32, Thm. 3.5.8] for unbalanced heights suggests it is of a very different nature.
@@ 9.7. Products
We next consider compatibility of our conjectures with respect to products. We focus on the balanced case as it is more fundamental. Let $G_{1}$ and $G_{2}$ be finite étale group schemes over $k$ with balanced heights $H_{1}$ and $H_{2}$. We define a height $H:=H_{1}\boxtimes H_{2}$ on $G:=G_{1}\times G_{2}$ by taking the product of height functions. To ensure that a product of balanced heights is balanced, we assume that the Fujita invariants are equal.
@@ 10.1. $S_{n}$-extensions
The orbifold line bundle $\Delta$ is balanced on $BS_{n}$.
@@
Therefore in this case the adjoint Iitaka fibration is trivial and the discriminant is a balanced height function, thus we are in the situation of Conjecture 9.1.
@@ 10.2.2. Discriminant
The subgroup generated by the minimal index elements is $M=\{(),(1,3),(2,4),(1,3)(2,4)\}$. The adjoint Iitaka fibration of the orbifold line bundle $\Delta$ is given by $BD_{4}\to BC_{2}$. In particular $\Delta$ is not balanced.
@@
The discriminant is not a balanced height function in this case. The adjoint Iitaka fibration is by definition the map which sends a $D_{4}$-extension $L/k$ to its quadratic subfield $L^{(1,3),(2,4)}$. By the Galois correspondence this is the unique quadratic subfield of the corresponding $D_{4}$-quartic $L^{(1,3)}$.
@@ 10.2.3. An Artin conductor
We have $a(C)=1,b(C)=2$. The orbifold line bundle $C$ is balanced.
@@ 10.3. Klüners’s counter-example
The group generated by these two conjugacy classes is $C_{3}\times C_{3}$ so the discriminant is not balanced and the Iitaka fibration is $BG\to BC_{2}$.
@@ 10.4. A new dihedral counter-example
In the case $n=3$ we obtain $S_{3}$, and one is counting non-cyclic cubic fields of bounded radical discriminant. The collection of points in the image of $BG_{3}$ is weakly accumulating: it needs to be removed before equidistribution is obtained. These correspond to exactly the pure cubic extensions; cf. Example 3.37. In fact this has been proven by Shankar and Thorne [84, Thm. 4].
@@ 10.5. $A_{4}$-quartics of bounded discriminant
There is in fact already a conjecture regarding this problem in [29, §2.7], but the expression is more complicated than ours and involves a limit process and sorting $A_{4}$-quartics by their cubic resolvent. It is not immediate how to compare the two conjectures directly. Nonetheless, in [29, §2.7] the numerical value $0.074\ldots$ is given for the leading constant, and a simple computation shows that our constant satisfies $2c(\mathbb{Q},A_{4},\Delta)=0.0729\ldots$ after considering the product over $0003<p<100,$ 000. (In private communication, Henri Cohen has confirmed that this minor discrepancy should not a problem and an artifact of using a least squares approximation in their calculation). The convergence is very slow, as one would expect, since it involves a conditionally convergent product; a faster converging product can be obtained by using the convergence factors from Theorem 8.17. We count isomorphism classes of fields, whereas [29] counts subfields of $\bar{\mathbb{Q}}$. To compare the two counts, we note that there are $4$ distinct embeddings of any $A_{4}$-quartic into $\bar{\mathbb{Q}}$, which matches with our factor $2$ in front and additional factor $1/2$ in the leading constant.
@@
with the latter two classes being swapped by $\operatorname{Gal}(\mathbb{Q}(\zeta_{3})/\mathbb{Q})$. It follows that $a(\Delta)=1/2$ and $b(\mathbb{Q},\Delta)=2$ (taking into account the Galois action). Since all conjugacy classes have minimal index, it follows that together they generate the group and hence the Iitaka fibration is trivial. Moreover the discriminant is the square of an anticanonical height; this may explain why this problem is so difficult, since the anticanonical height is generally the most difficult height in Manin’s conjecture.
We have $\widehat{A}_{4}=\mu_{3}$, so there are no non-trivial $1$-dimensional characters defined over $\mathbb{Q}$. The effective cone constant is thus $a(\Delta)^{b(\mathbb{Q},\Delta)-1}=1/2$. Since our height is a power of an anticanonical height, the relevant Brauer group is the unramified Brauer group.
@@
Finally Conjecture 9.3 states that we should remove $\widetilde{K}$ for which $\widetilde{K}\cap\mathbb{Q}(\mu_{6})\neq\mathbb{Q}$. However there are no such fields; indeed we have $\mathbb{Q}(\mu_{6})=\mathbb{Q}(\sqrt{-3})$, but $\widetilde{K}$ cannot contain a quadratic subfield as $A_{4}$ has no subgroup of index $2$.
@@ 10.6. $A_{4}$-quartics with a transcendental Brauer group
We now consider a different height function which exhibits new behaviour. Namely, there exists a balanced height function on $A_{4}$ for which the partially unramified transcendental Brauer group is non-trivial.
@@ 10.6.1. Numerical experiments
Let $\mathcal{C}$ consist of the minimal conjugacy classes $(1,2,3),(1,3,2)$. Note that we have $a(C)=1$ and $b(\mathbb{Q},C)=1$. The height is balanced since $(1,2,3)$ and its conjugate $(1,2,4)$ generate $A_{4}$.